OpenStudy (anonymous):

Is (8, -1) a solution of the system, 2x + 3y = 13 and x - y = 9?

OpenStudy (anonymous):

wt do u feel?

OpenStudy (anonymous):

I Say yes.

OpenStudy (amistre64):

lol ... yes is a good choice ;)

OpenStudy (anonymous):

then as u feel yes cross check it by pluggin in the value of x and just go ahead!!

OpenStudy (anonymous):

because NO is not good .. but the best one is "who cares ?" :P

OpenStudy (anonymous):

it is solution because if we put x=8 and y=-1 equation is satisfied

OpenStudy (anonymous):

is (-2, -5) a solution of the system, 3x + 2y = 4 and 2x - 5y = -29? So what would dis be?

OpenStudy (anonymous):

no because 3x+2y=4is not satisfied by (-2,-5)

OpenStudy (anonymous):

How are uu geting the answer thought?

OpenStudy (anonymous):

because if we put x=-2 and y=-4 in 3x+2y as 3(-2)+2(-5) we get -16 not 4 and the coordinate (-2,-5) must satisfy both equations of given system?

OpenStudy (anonymous):

oooooo

OpenStudy (anonymous):

yes now give me a medal and become my fan lolz

OpenStudy (anonymous):

thanks keona:)

OpenStudy (anonymous):

Youre Welcome Sumbul <3 :)!!

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

lol.